A north american bond pays semi-annual cooupons of 20$ and matures in 10 years. The current YTM is 5% and the face value is 1000$. What will be the price of this bond in 3 years, if the YTM in 3 years is 100 basis points higher than thw current YTM.


1
Expert's answer
2022-02-09T15:57:22-0500

100 basis points=1%


"YTM (3)=5+0.01\times5=5.05"


"P=\frac{40}{1+0.0505}+\frac{40}{1+0.0505}^2+\frac{1000}{1+0.0505}^3=936.93"